OpenStudy (anonymous):

4x-4 is greater than or equal to -3x^2 solve algebraically

OpenStudy (anonymous):

start with \[3x^2+4x-4\ge 0\]

OpenStudy (anonymous):

then see if you can factor to find the zeros

OpenStudy (anonymous):

OK

OpenStudy (anonymous):

I cant get it

OpenStudy (anonymous):

you get that last problem?

OpenStudy (anonymous):

No i didnt get that one either

OpenStudy (anonymous):

I need to find the solution set for this one

OpenStudy (anonymous):

\[3x^2+4x-4\geq 0\\(3x-2)(x+2)\geq 0\]

OpenStudy (anonymous):

zeros are at \(-2,\frac{2}{3}\) and since this is a parabola that faces up, it will be positive outside the zeros 

OpenStudy (anonymous):

two intervals \[(-\infty,-2]\cup [\frac{2}{3},\infty)\]

OpenStudy (anonymous):

so what is the solution set be

OpenStudy (anonymous):

then ones i wrote above

OpenStudy (anonymous):

what about use the remainder hteorem to find the remainder when f(x) is divided by x-1 than factor the theorem to determine whether x-1 is a factor of f(x)=2x^4-5x^3+4x-1

OpenStudy (anonymous):

remainder is \(f(1)\)

OpenStudy (anonymous):

i.e. \[2-5+4-1=0\]

OpenStudy (anonymous):

and since \(f(1)=0\) that makes \(x-1\) a factor of \[2x^4-5x^2+4x-1\]

OpenStudy (anonymous):

Do you have time to keep helping me?